OpenStudy (anonymous):

Find the value of the 10th term in the sequence of partial sums of 5(2/3)^n.

OpenStudy (tkhunny):

1: 5 2: 5+ 5(2/3) 3: 5+ 5(2/3) + 5(2/3)^2 = 5(1 + 2/3 + (2/3)^2) 4: 5(1 + 2/3 + (2/3)^2 + (2/3)^3) What else?

OpenStudy (anonymous):

I understand the concept, but I'm trying to compute it using the formula a(1-r^n)/(1-r) but am not sure which values go where :$

OpenStudy (tkhunny):

Right. What's preventing you from doing that? 3: 5+ 5(2/3) + 5(2/3)^2 = 5(1 + 2/3 + (2/3)^2) = \(5\cdot\dfrac{1 - (2/3)^3}{1-(2/3)}\) 4: 5(1 + 2/3 + (2/3)^2 + (2/3)^3) = \(5\cdot\dfrac{1 - (2/3)^4}{1-(2/3)}\)